# Identifying multi-compartment Hodgkin-Huxley models with high-density extracellular voltage recordings

### v1

@@ 4 Validation on simple morphologies
Figure 2: Validation on simple morphologies. Panels AC show results for a single-compartment cell. A: Estimated ${g}_{\mathrm{K}}$ over training iterations for three random initializations (mean +/- 1 sd). B: Distribution of inferred ${g}_{\mathrm{L}}$ under different observation noise levels $\sigma_{\text{obs}}$ (colored dashed line indicate the mean). C: Distribution of inferred ${g}_{\text{Na}}$ for $\sigma_{\text{obs}}=3\,\mu\text{V}$ for MSE minimization and EKF. Panels DF show results for a model with twelve compartments. D: A multi-compartment neuron and two sets of voltage recording grids (red and purple). The EKF correctly identifies the true location of the neuron (EKF in blue, true locations as black line). E: Recovery of membrane voltage in a compartment of the first branch, averaged over 40 datasets with $\sigma_{\text{obs}}=1\,\mu\text{V}$. F: Inferred conductances $g_{\mathrm{L}}$, ${g}_{\mathrm{K}}$ and ${g}_{\mathrm{Na}}$ of the first branch, across the 40 aforementioned datasets.
4.1 Validation on single-compartment neuron model
@@
We first validated our method on simple synthetic neuron morphologies. We considered a single-compartment neuron equipped with sodium and potassium Hodgkin–Huxley channels. The detailed experimental setup is provided in Appendix B.1. We initially assumed that the biophysical model was correctly specified and that the dynamics were fully deterministic. The extracellular probe had a single recording site placed $10\mu\mathrm{m}$ from the neuron, and the cell was stimulated with constant current injection. Our goal was to infer the maximum conductances, assuming the cell’s position was known. We simulated 40 extracellular voltage recordings corrupted by Gaussian noise with a standard deviation of $3\mu\mathrm{V}$, reflecting a high signal-to-noise scenario. Maximum conductances were initialized uniformly within biologically plausible ranges. EKF-based inference accurately recovered the true parameters across multiple random seeds, with low variance and negligible bias (fig. 2A), supporting the validity of our approach.
@@
Next, we assessed the effect of observation noise. Neuropixels Ultra probes exhibit RMS noise levels around $7.4\mu\mathrm{V}$ in saline and approximately $9.4\mu\mathrm{V}$ in vivo (Ye et al., 2024). Moreover, in practice we expect voltage traces to be corrupted by currents from other nearby neurons, which our model treats as noise as well. Fixing the cell–probe distance at $10\mu\mathrm{m}$, we simulated 40 noisy EAPs at three increasing noise levels, ranging from $3\mu V$ to $30\mu V$. As expected, we found that higher noise leads to higher variance in the parameter estimates (fig. 2B). Nevertheless, estimates remained near the ground truth even at $30\mu\mathrm{V}$ RMS, which represents a challenging, high-noise regime (fig. B.2).
@@
We also examined the effect of increasing the distance between the cell and the recording site. Since extracellular voltage scales proportionally with $1/r$, the signal-to-noise ratio declined with distance. In line with this intuition, we found that larger distance between cell and recording site leads to higher variance in the estimate of membrane conductances (fig. B.1B).
Finally, we compared our approach to a simpler alternative used in the original JAXLEY paper (Deistler et al., 2024). Instead of maximizing the EKF marginal likelihood approximation, we directly minimized the mean squared error between the observed extracellular potentials and those predicted by a model with deterministic dynamics. Since the model is well-specified, we expect this simpler alternative to also perform well in this setting. Using the same setup ($10\ \mu\mathrm{m}$ distance, $3\ \mu\mathrm{V}$ observation noise), we fit the same 40 traces using both methods. The resulting estimates were nearly indistinguishable (fig. 2C), suggesting that the local linearization used by EKF introduces negligible error. Additional figures can be found in Appendix B.1.
4.2 Validation on a multi-compartment neuron model
@@
We next evaluated our method in a more realistic multi-compartment setting. We again assumed that the biophysical model is correctly specified and the dynamics are fully deterministic. Using Jaxley, we simulated a synthetic neuron with three branches, each subdivided into four compartments, for a total of twelve compartments. Each compartment contained standard Hodgkin–Huxley ion channels, and the maximum conductances were allowed to vary independently across branches, yielding nine biophysical parameters to infer. The detailed experimental setup can be found in Appendix B.2. Unlike in the single-compartment case, we did not assume full knowledge of the cell’s location: while we treated the number of branches, their connectivity, lengths, and radii to be known, the 3D spatial locations of the branch endpoints were treated as unknown. This introduced 18 additional geometric parameters, bringing the total number of parameters to 27.
@@
A central question is whether these spatial locations are identifiable from extracellular recordings. If all recording sites lie in a single 2D plane—as is the case for a single Neuropixel probe—any mirror reflection of the neuron across that plane yields identical extracellular signals (formalization in Appendix C). To account for this unidentifiability, we simulated recordings from two separate probe positions, ensuring that the sets of recording sites did not lie in the same plane (fig. 2D, red and purple). We also ensured that the sites span the entire neuron to capture signal from all branches.
We fixed the observation noise to $1\mu\mathrm{V}$ and generated 40 sets of extracellular voltage recordings using this setup. To select an initialization, we first fixed a single set of recordings and sampled multiple random initializations from biologically plausible parameter ranges. We then chose the initialization that yielded the highest marginal likelihood after training. This initialization was held fixed and reused across all 40 simulated datasets. In this setup, our method consistently recovered the 3D locations of the center of the neuron’s compartments (fig. 2D, blue vs black), the membrane voltage (fig. 2E), and the membrane conductances (fig. 2F), demonstrating the robustness and accuracy of biophysical and geometric inference in multi-compartment neuron models with unknown locations and membrane conductances. Additional figures are provided in Appendix B.2.
@@ 4.3 Robustness to model misspecification
In experimental settings, the true underlying biophysical dynamics of a neuron are rarely known with certainty. To evaluate the robustness of our approach under such model misspecification, we conducted experiments where the ground truth and inference models differ in their ion channel composition and dynamics. The detailed experimental setup and additional figures can be found in Appendix B.3.
@@
In our first experiment, we generated synthetic data from a single-compartment neuron model containing both classical Hodgkin–Huxley (HH) channels and two additional conductances: an M-type potassium current and a calcium current. During inference, we deliberately used a simplified model containing only the HH channels, omitting the additional ones (fig. 3A).
@@
We next evaluated the effect of incorrect gating kinetics. Here, we retained the correct channel types in the inference model but perturbed the rate constants $\alpha(V)$ and $\beta(V)$ used in gating variable dynamics, simulating errors in channel parameter specification (fig. 3B). Once again, EKF-based inference provided significantly more accurate estimates of both membrane voltage and channel states compared to MSE minimization. These results suggest that our approach can tolerate moderate misspecification in gating dynamics. Together, these experiments demonstrate that EKF-based inference, combined with stochastic modeling of dynamics, confers robustness to common forms of model misspecification.
@@ 5 Application to real morphologies
Figure 4: Inference of retinal ganglion cell (RGC) properties given multi-electrode array (MEA) recordings. A: RGC cell geometry and electrode layout. Darker shaded MEA sites match the voltage traces shown in panel B. B: Simulated extracellular recordings at ten sites for noise level $\sigma_{\text{obs}}=10\,\mu\text{V}$. C: Estimation of $g_{\mathrm{K}}$ in the dendrites for 10 optimization runs under noise level $\sigma_{\text{obs}}=0.1\mu\text{V}$. D: Distribution of inferred $g_{\mathrm{L}}$ for $\sigma_{\text{obs}}\in\{0.1,10\}\,\mu\text{V}$. EKF identifies true parameters for sufficiently low values of noise. E: Recovered membrane voltage in the soma (left) and in a distal dendritic compartment (right) for both noise levels (mean +/- 1 sd). EKF matches the true voltage dynamics.
@@ 5.1 Application to 2D retinal ganglion cell
We applied our method to a morphologically detailed model of a retinal ganglion cell (RGC), replicating conditions typical of in vitro studies using planar multi-electrode arrays (MEAs) (Vilkhu et al., 2024). In these setups, retinal tissue is placed ganglion-side down on the MEA, enabling simultaneous current injection and extracellular voltage recording from hundreds of electrodes (fig. 4A). We adopted a reconstructed RGC morphology from Vilkhu et al. (2024), consisting of a soma and a complex dendritic arbor discretized into 122 compartments. The membrane model includes modified Hodgkin–Huxley sodium and potassium channels with non-standard kinetics. Following Vilkhu et al. (2024), we assumed shared maximum conductances across all dendritic compartments, yielding a total of six parameters to estimate. For details of the model, see Appendix D.
We here assumed the spatial location of the neuron to be known and constrained to a 2D plane, as is typical in retinal MEA experiments (Madugula et al., 2022; Sekirnjak et al., 2006; Vilkhu et al., 2024). The cell was stimulated via somatic current injection, and extracellular voltages were recorded using a simulated MEA with $30\mu\mathrm{m}$ inter-electrode spacing, matching the geometry of the original 512-electrode array. To reduce computational cost while maintaining full spatial coverage, we used a subset of 169 electrodes spanning the entire neuron. For each noise level ($0.1\mu\mathrm{V}$ and $10\mu\mathrm{V}$), we simulated ten extracellular action potentials (EAPs, fig. 4B).
Given the high dimensionality of the latent state space $(d=610)$, we used our diagonal EKF for efficient inference. We used multiple random initializations and selected the one that yielded the highest marginal likelihood. Our method accurately recovered the underlying conductances, exhibiting only minor bias (fig. 4C). While this bias increased with higher observation noise (fig. 4D), performance remained robust even at noise levels typical of MEA recordings (e.g., $10\mu\mathrm{V}$ (Multi Channel Systems MCS GmbH, 2020)). Notably, the inferred model was still able to faithfully reconstruct the true membrane voltage trace, highlighting the reliability of the approach under realistic noise conditions (fig. 4E). Additional figures are provided in Appendix D.
@@ 5.2 Application to CA1 hippocampal pyramidal cell
Finally, we applied our approach to a morphologically detailed model of a rat CA1 hippocampal pyramidal neuron, based on the C10 model introduced by Cutsuridis et al. (2010). This model has been used in prior computational studies to examine hippocampal circuit function and memory encoding, capturing the integration and propagation of signals across dendritic and axonal compartments (Cutsuridis et al., 2010; Cutsuridis and Taxidis, 2013). We recreated the full morphology of the C10 cell, including the soma, axon, basal dendrites, and apical dendrites (fig. 5A). The cell was equipped with region-specific Hodgkin–Huxley-type channels whose conductances and gating dynamics varied by subcellular location (Cutsuridis et al., 2010; Poirazi et al., 2003b; Migliore et al., 1999). This introduced 24 distinct maximum conductances to infer. The neuron was stimulated via somatic current injection, and extracellular traces were recorded using a simulated Neuropixel Ultra probe. Our virtual probe matched the true probe’s geometry, including the number of measurement sites and inter-site spacing.
We treated the spatial location of the neuron as unknown up to a rigid-body transformation, adding six parameters for rotation and translation and bringing the total number of parameters to 30. We initialized conductances with biologically plausible values for CA1 neurons. The detailed experimental setup and additional figures are provided in Appendix E. Following Carnevale and Hines (2006), the cell was discretized into $60$ compartments, resulting in a latent state of dimension $360$. Since this model size remained computationally tractable, we ran the full EKF rather than using the diagonal approximation. To avoid being trapped in a local maximum, we ran 100 independent restarts and retained the solution with the highest final marginal log-likelihood (fig. 5B).
Our method successfully recovers the neuron’s spatial configuration by estimating the 3D locations of compartment centers (fig. 5A, black vs. blue), and accurately reconstructs membrane voltage (fig. 5D). It also recovers some of the maximum conductances, though not all (fig. 5C). This outcome is consistent with a well-documented phenomenon in conductance-based neuron models: parameter degeneracy, where multiple combinations of channel conductances can yield virtually indistinguishable voltage dynamics (Prinz et al., 2004; Sterratt et al., 2023). Despite this, our method reliably recovers the neuron’s functional dynamics and spatial structure, demonstrating that biophysical and geometric inference from high-density extracellular recordings is feasible in realistic, large-scale neuron models.
@@ 6 Conclusions, Limitations, and Future Work
We presented an approach for identifying detailed biophysical neuron models from high-density extracellular voltage recordings. Our method recovers not only membrane conductances but also the spatial position of a neuron relative to the recording probe, solely from extracellular data. We presented an approach to state inference and parameter estimation using an Extended Kalman Filter (EKF), which provides robustness to model misspecification and accurate inference even when the assumed dynamics deviate from the true underlying system.
@@
One limitation of using an EKF to estimate model parameters is its computational cost: while we have developed a diagonal approximation to the filtering distribution, running an EKF is still computationally more costly than MSE minimization, especially for very large models. On top of this, any approximation made by the EKF (linearizing the dynamics, diagonal approximation) could bias parameter recovery. While we empirically found that our method could accurately recover parameters across a range of tasks, the approximations might have stronger effects on other models. Moreover it is common for biophysical models to exhibit parameter degeneracy: different combinations of maximum conductances can produce similar voltage dynamics, which can limit the identifiability of individual biophysical parameters, even when predictions match the data well.
Finally, while our results demonstrate the feasibility of biophysical inference from extracellular recordings at the single-cell level in idealized settings, extending this framework to real in vivo conditions remains a significant challenge. In particular, real recordings involve complex networks of interacting neurons. A natural next step is to move beyond isolated cells and model local neural populations surrounding the probe. Given knowledge of the targeted brain region, one could incorporate priors on network structure and cell types informed by anatomical and physiological studies, enabling more comprehensive modeling of in vivo extracellular signals.

### v2

@@ 4 Validation on simple morphologies
Figure 2: Validation on simple morphologies. Panels AC show results for a single-branch cell. A: Estimated ${g}_{\mathrm{K}}$ over training iterations for three random initializations (mean +/- 1 sd). B: Distribution of inferred ${g}_{\mathrm{L}}$ under different observation noise levels $\sigma_{\text{obs}}$ (colored dashed lines indicate the mean). C: Distribution of inferred ${g}_{\text{Na}}$ for $\sigma_{\text{obs}}=1\,\mu\text{V}$ for various inference methods (orange lines indicate the means). Panels DF show results for a model with twelve compartments. D: A multi-branch neuron and two sets of voltage recording electrodes (red and purple). The EKF correctly identifies the true location of the neuron (EKF in blue, true locations as black lines). E: Conductances inferred across 40 datasets simulated with $\sigma_{\text{obs}}=1\,\mu\text{V}$ via EKF for the 3rd branch of the cell. F: Recovery of membrane voltages for EKF and diagEKF in a compartment of the 3rd branch, averaged over the 40 aforementioned datasets.
4.1 Validation on single-branch neuron model
@@
We first validated our method on a simple synthetic neuron: a 10-compartment branch with sodium and potassium HH channels (Appendix E.1). We initially assumed that the biophysical model was correctly specified and that the dynamics were fully deterministic. We stimulated the cell with constant current and recorded extracellular voltages from a single electrode placed $5\,\mu\mathrm{m}$ from the soma. The goal was to infer maximum conductances assuming that the cell’s position was known. We simulated 40 recordings corrupted by Gaussian noise ($\sigma=1\,\mu\mathrm{V}$) and initialized conductances within biologically plausible ranges. EKF-based inference accurately recovered the true parameters across random seeds, with negligible bias and variance (fig.˜2A), supporting the validity of our approach.
@@
Next, we assessed the effect of observation noise. Neuropixels Ultra probes exhibit RMS noise levels around $7.4\mu\mathrm{V}$ in saline and approximately $9.4\mu\mathrm{V}$ in vivo [ye2024ultra]. Moreover, in practice we expect voltage traces to be corrupted by currents from other nearby neurons, which our model treats as noise as well. Fixing the cell–probe distance at $5\mu\mathrm{m}$, we simulated 40 noisy EAPs at three increasing noise levels, ranging from $1\mu V$ to $30\mu V$. As expected, we found that higher noise leads to higher variance in the parameter estimates (fig.˜2B). Nevertheless, estimates remained very close to the ground truth even at $30\mu\mathrm{V}$ RMS, which represents a challenging, high-noise regime (fig.˜E.3).
@@
We also examined the effect of increasing the distance between the cell and the recording site. Since extracellular voltage scales proportionally with $1/r$, the signal-to-noise ratio declined with distance. In line with this intuition, we found that larger distance between cell and recording site leads to higher variance in the estimate of membrane conductances (fig.˜E.1B).
Finally, we compared multiple inference approaches—standard EKF, diagonal EKF, block-diagonal EKF, and direct mean-squared-error (MSE) minimization as used in the original Jaxley paper [deistler2024differentiable]. The MSE baseline replaces the probabilistic objective with deterministic dynamics and directly minimizes the squared error between observed and predicted extracellular potentials. Using the same setup ($5\ \mu\mathrm{m}$ distance, $1\ \mu\mathrm{V}$ observation noise) and fitting the same 40 random traces, we found that the standard EKF, block-diagonal EKF, and MSE approaches achieved nearly identical estimates (fig.˜2C). This confirms that the local linearization in the EKF introduces negligible bias and that the block-diagonal approximation provides an excellent surrogate for the full covariance in HH models. The diagonal EKF exhibited slightly higher bias but its voltage reconstructions remain highly accurate (fig.˜E.2). This is consistent with our analysis in Appendix B: only the first compartment lies near the electrode, and it has a single neighboring compartment, leading to minimal cross-visibility and making a diagonal covariance approximation efficient.
4.2 Validation on a multi-branch neuron model
@@
We next evaluated our method on a more realistic multi-branch setting, assuming a correctly specified biophysical model and fully deterministic dynamics. Using Jaxley, we simulated a synthetic neuron with three branches, each divided into four compartments (twelve in total), containing standard HH ion channels. The maximum conductances varied independently across branches, yielding nine biophysical parameters to infer. Details of the setup are provided in Appendix E.2. Unlike the single-branch case, we did not assume full knowledge of the cell’s spatial configuration: while branch number, connectivity, lengths, and radii were known, the 3D coordinates of branch endpoints were treated as unknown, introducing 18 additional geometric parameters for a total of 27.
@@
A central question is whether these spatial locations are identifiable from extracellular recordings. With electrodes confined to a single plane (e.g., a Neuropixel probe), mirror reflections of the neuron across the plane produce identical extracellular voltages (cf. Appendix D). To account for this unidentifiability, we simulated recordings from two separate probe positions, ensuring that the sets of recording sites did not lie in the same plane (fig.˜2D, red and purple). We also ensured that the sites span the entire neuron to capture signal from all branches.
We fixed the observation noise to $1\mu\mathrm{V}$ and generated 40 sets of extracellular voltage recordings. To initialize parameters, we optimized from multiple random starting points on one dataset and retained the best solution (highest marginal log-likelihood), which we then reused across all runs. In this setting, the EKF consistently recovered the 3D compartment locations (fig.˜2D, blue vs black), conductances (fig.˜2E), and membrane voltages (fig.˜2F), demonstrating the robustness and accuracy of biophysical and geometric inference in multi-compartment neuron models with unknown locations and membrane conductances. The diagonal EKF exhibited somewhat larger bias in the recovered conductances compared with the single-branch case (fig.˜E.6), but the voltage traces remained accurate up to small shifts (fig.˜2F), and the compartment positions were inferred correctly (fig.˜E.8). In this multi-branch geometry, compartments are long enough that most electrodes are close to only a single compartment, and each compartment has very few neighbors, which together limit cross-compartment measurement interactions and explains why the diagonal approximation still yields informative reconstructions (cf. Appendix B).
@@ 4.3 Robustness to model misspecification
In experimental settings, the true underlying biophysical dynamics of a neuron are rarely known with certainty. To evaluate the robustness of our approach under such model misspecification, we conducted experiments where the ground truth and inference models differ in their ion channel composition and dynamics. The detailed experimental setup and additional figures can be found in Appendix E.3.
@@
In our first experiment, we generated synthetic data from a single-branch cell containing both classical HH channels and two additional conductances: an M-type potassium current and a calcium current. During inference, we deliberately used a simplified model containing only the HH channels, omitting the additional ones (fig.˜3A).
@@
We next evaluated the effect of incorrect gating kinetics. Here, we retained the correct channel types in the inference model but perturbed the rate constants $\alpha(V)$ and $\beta(V)$ used in gating variable dynamics, simulating errors in channel parameter specification (fig.˜3B). Once again, EKF-based inference provided significantly more accurate estimates of both membrane voltage and channel states compared to MSE minimization. These results suggest that our approach can tolerate moderate misspecification in gating dynamics. Together, these experiments demonstrate that EKF-based inference, combined with stochastic modeling of dynamics, confers robustness to common forms of model misspecification.
@@ 5 Application to real morphologies
Figure 4: Inference of retinal ganglion cell (RGC) properties given multi-electrode array (MEA) recordings. A: RGC cell geometry and electrode layout. Darker shaded MEA sites match the voltage traces shown in panel B. B: Simulated extracellular recordings at ten sites for noise level $\sigma_{\text{obs}}=0.5\,\mu\text{V}$. C: Marginal log-likelihood during optimization for diagonal EKF. D: The heatmap of the voltage block of $\mathbf{H}^{\top}\mathbf{H}$ (log scale) shows that the diagonal approximation is expected to fail. E: Marginal log-likelihood for block-diagonal EKF. F: Fitted voltage traces in the soma. The block-diagonal EKF infers the voltages accurately while the diagonal EKF fails.
@@ 5.1 Application to 2D retinal ganglion cell
We applied our method to a morphologically detailed model of a retinal ganglion cell (RGC), replicating conditions typical of in vitro studies using planar multi-electrode arrays (MEAs) [vilkhu2024understanding]. In these setups, retinal tissue is placed ganglion-side down on the MEA, enabling simultaneous current injection and extracellular voltage recording from hundreds of electrodes (fig.˜4A). We adopted a reconstructed RGC morphology from vilkhu2024understanding, consisting of a soma and a complex dendritic arbor discretized into 122 compartments. The membrane model includes modified HH sodium and potassium channels with non-standard kinetics. We assumed shared maximum conductances across all dendritic compartments, yielding a total of six parameters to estimate [vilkhu2024understanding]. For details of the model, see Appendix F.
Here, we assumed the spatial location of the neuron to be known and constrained to a 2D plane, as is typical in retinal MEA experiments [madugula2022focal, sekirnjak2006electrical, vilkhu2024understanding]. Extracellular voltages were recorded using a simulated MEA with $30\mu\mathrm{m}$ pitch, matching the geometry of the original 512-electrode array. To reduce computation, we used a subset of 10 electrodes spanning the entire neuron.
Given the high dimensionality of the latent state space ($d=610$), we used our scalable EKF variants for efficient inference. Notably, the diagonal EKF performed poorly: assuming diagonal covariances, the marginal log-likelihood diverged from its value at the true parameters, indicating severe model mismatch (fig.˜4C). The reconstructed voltage traces were also inaccurate, failing to reproduce the true dynamics (fig.˜4F). This is precisely the failure mode predicted in Section 3.2: since each electrode records signal from many close compartments, $\mathbf{H}^{\top}\mathbf{H}$ cannot be diagonal (fig.˜4D). In contrast, the block-diagonal EKF converged to the true marginal likelihood (fig.˜4E) and accurately reconstructed both the membrane voltage traces (fig.˜4F) and the underlying conductances (fig.˜F.1).
@@ 5.2 Application to CA1 hippocampal pyramidal cell
Finally, we applied our approach to a morphologically detailed model of a rat CA1 hippocampal pyramidal neuron, based on the C10 model introduced by cutsuridis2010encoding. This model has been used in prior computational studies to examine hippocampal circuit function and memory encoding, capturing the integration and propagation of signals across dendritic and axonal compartments [cutsuridis2010encoding, cutsuridis2013deciphering]. We recreated the full morphology of the C10 cell, including the soma, axon, basal dendrites, and apical dendrites (fig.˜5A). The cell contained region-specific HH-type channels with subcellularly varying conductances and gating kinetics [cutsuridis2010encoding, poirazi2003arithmetic, migliore1999role], yielding 24 maximum conductances to infer. Extracellular signals were recorded with a simulated Neuropixel Ultra probe replicating the real probe’s geometry. To account for unknown spatial position, we added six rigid-body parameters (rotation and translation), for a total of 30 parameters. Further experimental details and figures are provided in Appendix G. Following carnevale2006neuron, the cell was discretized into $60$ compartments, resulting in a latent state of dimension $360$. We ran the block-diagonal EKF with over 100 restarts and retained the solution with the highest final marginal log-likelihood (fig.˜5B).
Our method successfully recovers the neuron’s spatial configuration (fig.˜5A, black vs. blue), and accurately reconstructs membrane voltage (fig.˜5D). It also recovers most of the maximum conductances, though not all (fig.˜5C). This outcome is consistent with a well-documented phenomenon in conductance-based neuron models: parameter degeneracy, where multiple combinations of channel conductances can yield virtually indistinguishable voltage dynamics [prinz2004similar, sterratt2023principles]. Despite this, our method reliably recovers the neuron’s functional dynamics and spatial structure, demonstrating that biophysical and geometric inference from high-density extracellular recordings is feasible in realistic, large-scale neuron models.
@@ 6 Conclusions, Limitations, and Future Work
We presented an approach for identifying detailed biophysical neuron models from high-density extracellular voltage recordings. Our method recovers not only membrane conductances but also the spatial position of a neuron relative to the recording probe, solely from extracellular data. We further developed a state-inference and parameter-estimation framework based on an Extended Kalman Filter (EKF), which provides robustness to model misspecification and accurate inference even when the assumed dynamics deviate from the true underlying system.
@@
A limitation of the EKF for parameter estimation is its computational cost: even with diagonal or block-diagonal approximations, it remains more expensive than MSE minimization, particularly for large models. Additionally, EKF approximations may bias parameter recovery: while the block-diagonal EKF recovered parameters accurately in our evaluations, the diagonal approximation was less reliable in certain cell–probe geometries. Moreover, biophysical models often exhibit parameter degeneracy, where different conductance combinations yield similar voltage dynamics, limiting identifiability even when predictions fit the data well.
Finally, while our results demonstrate the feasibility of biophysical inference from extracellular recordings at the single-cell level in idealized settings, extending this framework to real in vivo conditions remains a significant challenge. Real recordings involve complex networks of interacting neurons. A natural next step is to move beyond isolated cells and model local neural populations surrounding the probe. Given knowledge of the targeted brain region, one could incorporate priors on network structure and cell types informed by anatomical and physiological studies, enabling more comprehensive modeling of in vivo extracellular signals.